OpenStudy (kaliyah0524):

Solve for x: 1 over 3 (2x − 8) = 4.

OpenStudy (anonymous):

have ou tried working it out yet

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }(2x - 8)=4\]\[(2x - 8) = 4*3\]\[2x - 8 + 8 = 12 + 8\]\[2x = 20\] x = 10

OpenStudy (anonymous):

your not supposed to give the answer but yea that's right